# Parallel Exploration of the Optoelectronic Properties of (Sb,Bi)(S,Se)(Br,I) Chalcohalides

### v1

@@ I Introduction
Lead-halide perovskites have transformed the field of optoelectronics due to their remarkable material properties [1, 2, 3, 4, 5], including high absorption coefficients [6], long carrier diffusion lengths [7], and an unusual tolerance to defects [8]. Despite these advantages, concerns about lead toxicity [9, 10] and long-term instability [11, 12, 13] have catalyzed the search for alternative materials. Chalcogenides, in contrast, have been explored as another promising class of optoelectronic materials, offering excellent chemical and thermal stability, but they generally fall short of matching the superior optoelectronic performance of halide perovskites [14]. The ultimate goal in developing next-generation optoelectronic technologies is to discover a new family of materials that combines the defect tolerance and electronic properties of halide perovskites with the exceptional environmental and operational stability of chalcogenides.
@@
In this work, we investigate the optoelectronic properties of a combinatorial subset of heavy pnictogen-based ternary chalcohalides with the general formula (Sb,Bi)(S,Se)(Br,I). These eight compounds all crystallize in the same quasi-one-dimensional orthorhombic structure, providing an ideal platform for comparative studies. We present a two-step synthesis route, in which chalcogenide precursors are first formed via thermal co-evaporation, followed by high-pressure reactive annealing to incorporate the halogen and form the final ternary phase. We first characterize their structural and morphological properties, alongside optical measurements to evaluate light absorption and emission behavior. Based on these results, we identify the most promising compounds for more in-depth photoluminescence (PL) studies, including power- and temperature-dependent measurements, as well as excitation intensity-dependent time-resolved PL spectroscopy (TRPL). These measurements are used to assess the nature of the emissive states, charge-carrier recombination dynamics, and electron-phonon interactions, offering insight into the fundamental processes that govern the intrinsic optoelectronic performance in this class of materials.
@@ II Methods
Following deposition, the chalcogenide precursor was placed in a Petri dish along with the corresponding halide source – SbI3 (Thermo-Scientific, 99.999%), SbBr3 (Thermo-Scientific, 99.995%), BiI3 (Puratronic, 99.999%), or BiBr3 (Thermo-Scientific, 99%). For example, Sb2Se3 was combined with SbI3 to form SbSeI. The Petri dish was then sealed within a steel tubular furnace under an inert Ar atmosphere and subjected to high-temperature, high-pressure annealing. The annealing conditions were optimized for each chalcohalide composition, as detailed in Table 1. All purities specified are metals basis.
@@ III RESULTS
First, we characterized the structural properties of the chalcohalide samples synthesized on glass/Mo substrates. We quantified the chemical compositions using EDS mapping, examined microstructural morphologies via top-view SEM imaging, and determined the crystal structures from XRD patterns. The identification of secondary phases is particularly important, as the assessment of intrinsic material properties of a given compound is only meaningful when based on phase-pure samples.
@@
Se-based compounds typically exhibit lower PL yields than their S-based counterparts due to a combination of intrinsic electronic and structural factors. The larger atomic radius and weaker metal–selenium bonds in Se compounds tend to increase the density of structural defects and non-radiative recombination centers, both of which can suppress radiative emission. Recent work by López et al. [70] demonstrated that Se-based pnictogen chalcohalides (e.g., BiSeI and BiSeBr) exhibit higher equilibrium concentrations of native defects – particularly selenium vacancies – than their sulfur-based analogues (e.g., BiSI and BiSBr). This trend is attributed to the lower formation energies and higher volatility of selenium under typical synthesis conditions. Furthermore, the reduced bandgaps of Se-based compounds favor thermal population of non-radiative states, which can enhance phonon-assisted decays [71, 72]. The stronger spin-orbit coupling in selenium, due to its higher atomic mass, can also facilitate exciton relaxation into dark states, further reducing radiative efficiency. These effects have been widely observed in other systems. For instance, monolayer MoSe2 exhibits significantly lower PL intensity than MoS2 under identical conditions, primarily due to increased non-radiative decay pathways and dark exciton formation [73]. Similarly, CdSe quantum dots without core-shell passivation generally show lower PL quantum yields than CdS counterparts [74], highlighting a broader trend of diminished PL efficiency in Se-based semiconductors.
@@
To quantify the temperature-dependence of the PL emission, we fit the peak at each temperature with a single Gaussian and track the extracted peak position and spectral width. These trends are shown in Figures 4(e) and 4(f), respectively. As both the thermal shift of the bandgap and the broadening of the emission peak are typically attributed to electron-phonon interactions, we model them using Bose-Einstein statistics. The temperature dependence of the bandgap is described by:
@@
In contrast, the fitted parameters for BiSeI show a significantly higher phonon coupling strength ( meV) and an unusually large characteristic phonon energy ( meV), well beyond the typical energy range of lattice vibrations. This suggests that while the Bose–Einstein model generally describes the temperature dependence of PL in BiSeI, it likely oversimplifies the true broadening mechanisms, potentially involving anharmonic lattice dynamics or multiple phonon modes. These anomalies align with the recent insights into the native defect chemistry of this materials system from López et al. [70], who showed that the equilibrium concentration of selenium vacancies (VSe) in BiSeI is relatively high due to a low formation enthalpy under Se-poor conditions. These vacancies introduce deep charge transition states that drive non-radiative recombination, significantly reducing photocarrier lifetimes and radiative efficiency. In contrast, BiSeBr benefits from higher defect formation enthalpies – particularly for antisite defects – due to larger ionic radius mismatches, which effectively suppress defect formation. Collectively, these results illustrate that the stronger apparent phonon coupling and enhanced PL broadening observed for BiSeI not only reflect more intense electron–phonon interactions but also highlight the severe impact of high defect densities on non-radiative carrier decay.
@@ IV Discussion
In addition to defect chemistry, vibrational effects also play a central role in the optoelectronic performance of these chalcohalides. Figure 5 shows the projected phonon density of states (DOS) for BiSI, calculated using density functional theory (DFT), revealing a pronounced phonon gap just below the sulfur-related phonon band (25–30 meV), which closely aligns with the experimentally extracted characteristic phonon energy . This gap originates from the large atomic mass differences between the heavy pnictogen and the much lighter sulfur atoms and is consistently observed across all sulfide-based compounds (Figure S3 in the Supplementary Information). While this gap can reduce phonon scattering pathways —- potentially improving radiative efficiency – it may also enhance energy dissipation through specific high-energy modes, promoting non-radiative recombination. In contrast, the corresponding selenium-based compounds lack this phonon gap, as selenium-related phonon modes extend into the previously forbidden vibrational range. This filled gap intensifies carrier–phonon scattering, providing a compelling vibrational explanation for the broader PL observed in BiSeI.
@@ Acknowledgements
The work presented here is supported by the Carlsberg Foundation, grant CF24-0200. C.L. acknowledges support from the Spanish Ministry of Science, Innovation and Universities under an FPU grant. C.C. acknowledges support by MICIN/AEI/10.13039/501100011033 and ERDF/EU under the grants PID2023-146623NB-I00 and PID2023-147469NB-C21 and by the Generalitat de Catalunya under the grants 2021SGR-00343, 2021SGR-01519 and 2021SGR-01411. Computational support was provided by the Red Española de Supercomputación under the grants FI-2024-1-0005, FI-2024-2-0003, FI-2024-3-0004, FI-2024-1-0025, FI-2024-2-0006, FI-2025-1-0015 and FI-2025-1-0001.
@@
The data that support the findings of this study are available from the corresponding author upon request.

### v2

@@ I Introduction
Lead-halide perovskites have transformed the field of optoelectronics due to their remarkable material properties [1, 2, 3, 4, 5], including high absorption coefficients [6], long carrier diffusion lengths [7], and an unusual tolerance to defects [8]. Despite these advantages, concerns about lead toxicity [9, 10] and long-term instability [11, 12, 13] have catalyzed the search for alternative materials. Chalcogenides, in contrast, have been explored as another promising class of optoelectronic materials, offering excellent chemical and thermal stability, but they generally fall short of matching the superior optoelectronic performance of halide perovskites [14, 15]. The ultimate goal in developing next-generation optoelectronic technologies is to discover a new family of materials that combines the defect tolerance and electronic properties of halide perovskites with the exceptional environmental and operational stability of chalcogenides.
@@
In this work, we investigate the optoelectronic properties of a combinatorial subset of heavy pnictogen-based ternary chalcohalides with the general formula (Sb,Bi)(S,Se)(Br,I). These eight compounds all crystallize in the same quasi-one-dimensional orthorhombic structure with space group Pnma, providing an ideal platform for comparative studies. We present a two-step synthesis route, in which chalcogenide precursors are first formed via thermal co-evaporation, followed by high-pressure reactive annealing to incorporate the halogen and form the final ternary phase. We first characterize their structural and morphological properties, alongside optical measurements to evaluate light absorption and emission behavior. Based on these results, we identify the most promising compounds for more in-depth photoluminescence (PL) studies, including power- and temperature-dependent measurements, as well as excitation intensity-dependent time-resolved PL spectroscopy (TRPL). These measurements are used to assess the nature of the emissive states, charge-carrier recombination dynamics, and electron-phonon interactions, offering insight into the fundamental processes that govern the intrinsic optoelectronic performance in this class of materials.
@@ II Methods
Following deposition, the chalcogenide precursor was placed in a Petri dish along with the corresponding halide source – SbI3 (Thermo-Scientific, 99.999%), SbBr3 (Thermo-Scientific, 99.995%), BiI3 (Puratronic, 99.999%), or BiBr3 (Thermo-Scientific, 99%). For example, Sb2Se3 was combined with SbI3 to form SbSeI. The Petri dish was then sealed within a steel tubular furnace under an inert Ar atmosphere and subjected to high-temperature, high-pressure annealing. The annealing conditions were optimized for each chalcohalide composition, as detailed in Table 1. All purities specified are metals basis. The synthesis process has been schematically illustrated in Figure S1. See Ref [32] for further synthesis details.
@@ III RESULTS
First, we characterized the structural properties of the chalcohalide samples synthesized on glass/Mo substrates. We quantified the chemical compositions using EDX mappings, examined microstructural morphologies via top-view SEM imaging, and determined the crystal structures from XRD patterns. The identification of secondary phases is particularly important, as the assessment of intrinsic material properties of a given compound is only meaningful when based on phase-pure samples.
@@
Se-based compounds typically exhibit lower PL yields than their S-based counterparts due to a combination of intrinsic electronic and structural factors. The larger atomic radius and weaker metal–selenium bonds in Se compounds tend to increase the density of structural defects and non-radiative recombination centers, both of which can suppress radiative emission. Recent work by López et al. [72] demonstrated that Se-based pnictogen chalcohalides (e.g., BiSeI and BiSeBr) exhibit higher equilibrium concentrations of native defects – particularly selenium vacancies – than their sulfur-based analogues (e.g., BiSI and BiSBr). This trend is attributed to the lower formation energies and high volatility of selenium under typical synthesis conditions. Furthermore, the reduced bandgaps of Se-based compounds favor thermal population of non-radiative states, which can enhance phonon-assisted decays [73, 74]. The stronger spin-orbit coupling in selenium, due to its higher atomic mass, can also facilitate exciton relaxation into dark states, further reducing radiative efficiency. These effects have been widely observed in other systems. For instance, monolayer MoSe2 exhibits significantly lower PL intensity than MoS2 under identical conditions, primarily due to increased non-radiative decay pathways and dark exciton formation [75]. Similarly, CdSe quantum dots without core-shell passivation generally show lower PL quantum yields than CdS counterparts [76], highlighting a broader trend of diminished PL efficiency in Se-based semiconductors.
@@
To quantify the temperature-dependence of the PL emission, we fit the peak at each temperature with a single Gaussian () and track the extracted peak position and spectral width. These trends are shown in Figures 4(e) and 4(f), respectively. As both the thermal shift of the bandgap and the broadening of the emission peak are typically attributed to electron-phonon interactions, we model them using Bose-Einstein statistics. The temperature dependence of the bandgap is described by:
@@
In contrast, the fitted parameters for BiSeI show a significantly higher phonon coupling strength ( meV) and an unusually large characteristic phonon energy ( meV), well beyond the typical energy range of lattice vibrations. This suggests that while the Bose–Einstein model generally describes the temperature dependence of PL in BiSeI, it likely oversimplifies the true broadening mechanisms, potentially involving anharmonic lattice dynamics or multiple phonon modes. These anomalies align with the recent insights into the native defect chemistry of this materials system from López et al. [72], who showed that the equilibrium concentration of selenium vacancies (VSe) in BiSeI is relatively high due to a low formation enthalpy under Se-poor conditions. These vacancies introduce deep charge transition states that drive non-radiative recombination, significantly reducing photocarrier lifetimes and radiative efficiency. Although the capture coefficients of these defects are calculated to be modest, their high concentrations can substantially limit device performance. The detrimental impact of these vacancies may be mitigated through cation-poor synthesis conditions and strategic anion substitutions, offering potential pathways to improve the optoelectronic performance of BiSeI. In contrast, BiSeBr benefits from higher defect formation enthalpies – particularly for antisite defects – due to larger ionic radius mismatches, which effectively suppress defect formation. Collectively, these results illustrate that the stronger apparent phonon coupling and enhanced PL broadening observed for BiSeI not only reflect more intense electron–phonon interactions but also highlight the severe impact of high defect densities on non-radiative carrier decay.
@@ IV Discussion
In addition to defect chemistry, vibrational effects also play a central role in the optoelectronic performance of these chalcohalides. Figure 5 shows the total phonon density of states (DOS) for BiSI, calculated using density functional theory (DFT), revealing a pronounced phonon gap just below the sulfur-related phonon band (25–30 meV), which closely aligns with the experimentally extracted characteristic phonon energy . This gap originates from the large atomic mass differences between the heavy pnictogen and the much lighter sulfur atoms and is consistently observed across all sulfide-based compounds (Figure S4 in the Supplementary Information). While this gap can reduce phonon scattering pathways – potentially improving radiative efficiency – it may also enhance energy dissipation through specific high-energy modes, promoting non-radiative recombination. In contrast, the corresponding selenium-based compounds lack this phonon gap, as selenium-related phonon modes extend into the previously forbidden vibrational range. This filled gap intensifies carrier–phonon scattering, providing a compelling vibrational explanation for the anomalous electron-phonon coupling strength observed in BiSeI.
@@ Acknowledgements
The work presented here is supported by the Carlsberg Foundation, grant CF24-0200. This project received funding from the European Union’s Horizon research and innovation programme under grant agreements number 866018 (SENSATE); by the Spanish Ministry of Science and Innovation project number PCI2023-145971-2 (ACT-FAST) from the CETP-Partnership Program 2022; and by the Cost Association project number CA-706 21148 (Renew-PV). This work is also part of Maria de Maeztu Units of Excellence Programme CEX2023-001300-M, funded by MICIU/AEI/10.13039/501100011033. The authors from UPC and IREC belong to the Micro and Nanotechnologies for Solar Energy 709 Group (MNTSolar) Consolidated Research Group of the “Generalitat de Catalunya” (2021 SGR 01286). E.S. acknowledges the ICREA Academia program. I.C., acknowledges Inno-PV (PID2022-140226OB-C31, PID2022-140226OB-C32), funded by MCIN/AEI/10.13039/501100 011033/FEDER. C.L. acknowledges support from the Spanish Ministry of Science, Innovation and Universities under an FPU grant. C.C. acknowledges support by MICIN/AEI/10.13039/501100011033 and ERDF/EU under the grants PID2023-146623NB-I00 and PID2023-147469NB-C21 and by the Generalitat de Catalunya under the grants 2021SGR-00343, 2021SGR-01519 and 2021SGR-01411. Computational support was provided by the Red Española de Supercomputación under the grants FI-2024-1-0005, FI-2024-2-0003, FI-2024-3-0004, FI-2024-1-0025, FI-2024-2-0006, FI-2025-1-0015 and FI-2025-1-0001.
@@
All relevant data and code for analysis generated in the course of this work is freely available at https://doi.org/10.5281/zenodo.15614394.